OpenStudy (anonymous):

y^2 = x . is this the equation for a parabola?

OpenStudy (anonymous):

equation of parabola is in y= second degree x. in this case , it is y^2= x. taking sqrt on both side you get y=sqrt(x) and y= -sqrt(x).

OpenStudy (anonymous):

No its y= x^2...

OpenStudy (anonymous):

so if you are in highschool, i would go with a no, although y^&2= x is basically a parabola around the x axis.

OpenStudy (anonymous):

yeah. that's what I was wondering because I thought a parabola had to be a function which this equation is not.